I need help with these questions ASAP

34&35

Note: It is recommended that you save your response as you complete each question. Question 35 (2.5 points) Panthers, Inc. has a capital structure of 20% debt and 80% equity. The tax rate is 40%. The firms bonds currently trade in the market for $950. These face value $1,000 bonds have a coupon rate of 6%, paid semiannually, with 10 years to maturity. The firms common stock trades for $20 per share. The firm just paid a dividend of $2. Future dividends are expected to grow at 3% per year forever. Panthers WACC is 96. 1) 7.781 2) 9.089 3) 11.443 4) 13.528 5) 15.170 Save Page 50 Pr
media%2F6a4%2F6a4d58b8-eeee-4579-8d48-21
Answer #1

35) Cost of debt can be calculated using I/Y function on a calculator

N = 10 x 2 = 20, PMT = 6% x 1000 / 2 = 30, PV = -950, FV = 1000 => Compute I/Y = 3.35% (semi-annual)

Annual Cost of debt, rd = 2 x 3.35% = 6.69%

Cost of equity, re = D0 x (1 + g) / P + g = 2 x (1 + 3%) / 20 + 3% = 13.30%

WACC = wd x rd x (1 - tax) + we x re = 20% x 6.69% x (1 - 40%) + 80% x 13.30% = 11.443%

34) 3 is correct. Choose Project B as it has higher IRR (and more importantly, higher NPV) than Project A.